# Elliptic boundary-value problems in some distribution spaces of generalized smoothness

### v1

@@ 3. Applications to elliptic problems
Assertion (i) in the $\ell=0$ case follows from [31, Theorem 5.3] with regard for [9, Theorem 3.4] (see also [29, Theorem 1], where the proof is given). The $\ell\geq 1$ case is reduced to the previous one because the differential operator $\partial/\partial x_{j}$ acts continuously from $H^{s,\varphi}_{p}(\mathbb{R}^{n})$ to $H^{s-1,\varphi}_{p}(\mathbb{R}^{n})$ for every $s\in\mathbb{R}$. Assertion (ii) is proved by a modification of the reasoning given in [64, Proof of Theorem 13.2] in the $\ell=0$ case. Namely, we may assume that $0\in V$; then (3.23) implies the continuous embedding $H^{\ell+n/p,\varphi}_{p}(V_{0})\hookrightarrow C^{\ell}(\overline{V_{0}})$ for a certain ball $V_{0}$ centered at 00, which gives the estimate
@@
for every $w\in\mathcal{S}(\mathbb{R}^{n})$ and each partial derivative $\partial^{\alpha}$ of the order $|\alpha|\leq\ell$. Hence, the delta-function $\delta$ and all its partial derivatives $\partial^{\alpha}\delta$ with $|\alpha|\leq\ell$ belong to the dual of $H^{\ell+n/p,\varphi}_{p}(\mathbb{R}^{n})$. Since this dual equals $H^{-\ell-n/p,1/\varphi}_{r}(\mathbb{R}^{n})$ [64, Section 13, Subsection 3], we have the inclusion $\partial^{\alpha}\Phi\delta=\Phi(\partial^{\alpha}\delta)\in H^{-\ell-n/p}_{r}(\mathbb{R}^{n})$ whatever $0\leq|\alpha|\leq\ell$, with $\Phi w:=\mathcal{F}^{-1}[\varphi^{-1}(\langle\xi\rangle)\mathcal{F}w(\xi)]$ whenever $w\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$. Hence, $\Phi\delta\in H^{-n/p}_{r}(\mathbb{R}^{n})$ (see, e.g., [61, Theorem 2.3.8(ii)]), which gives $\delta\in H^{-n/p,1/\varphi}_{r}(\mathbb{R}^{n})$. The last inclusion means that the function $\mu(\xi):=\langle\xi\rangle^{n/p}\varphi(\langle\xi\rangle)$ of $\xi\in\mathbb{R}^{n}$ satisfies $\mathcal{F}^{-1}\mu^{-1}(\xi)\in L_{r}(\mathbb{R}^{n})$. This inclusion is equivalent to (3.22) due to [64, Theorem 13.2] and [31, Theorem 5.3]. Thus, (3.23) implies (3.22).
@@ 4. Applications to parameter-elliptic boundary-value problems
    For arbitrarily fixed point $x\in\Gamma$, vector $\xi\in\nobreak\mathbb{R}^{n}$, tangent to the boundary $\Gamma$ at the point $x$, and the parameter $\lambda\in K$ such that $|\xi|+|\lambda|\neq 0$, the polynomials $B^{\circ}_{j}(x;\xi+\tau\nu(x),\lambda)$, $j=1,\ldots,q$, in $\tau$ are linearly independent modulo the polynomial $\prod_{j=1}^{q}(\tau-\tau^{+}_{j}(x;\xi,\lambda))$. Here, $\tau^{+}_{1}(x;\xi,\lambda),\ldots,\tau^{+}_{q}(x;\xi,\lambda)$ are all $\tau$-roots of the polynomial $A^{\circ}(x;\xi+\tau\nu(x),\lambda)$ that have the positive imaginary part and are written with regard for their multiplicity.
We have used the following designations:
@@
    There exists a number $\lambda_{0}>0$ such that, for each $\lambda\in K$ with $|\lambda|\geq\nobreak\lambda_{0}$, operator (4.3) is an isomorphism of $H^{s,\varphi}_{p}(\Omega)$ onto $\Xi^{s,\varphi}_{p}(\Omega,\Gamma)$ whatever $p\in(1,\infty)$, $s>m+1/p$, and $\varphi\in\Upsilon$.
@@
    $\quad+\sum_{j=1}^{q}\,\bigl{(}\|B_{j}(\lambda)u,B^{s-m_{j}-1/p,\varphi}_{p}(\Gamma)\|+|\lambda|^{s-m_{j}-1/p}\varphi(|\lambda|)\,\|B_{j}(\lambda)u,B^{0}_{p}(\Gamma)\|\bigr{)}\leq$
@@
Note that $L_{p}(\Omega)=H^{0}_{p}(\Omega)$ with equality of norms. Since the ellipticity of the parameter-dependent problem (4.1), (4.2) in $K$ implies its ellipticity in the angle $-K:=\{-\lambda:\lambda\in K\}$, the conclusions of this theorem is also valid for $\lambda\in-K$ with $|\lambda|\gg 1$.
@@
$\begin{split}\|\mathcal{R}f,B^{s,\varphi}_{p}(\mathbb{R}^{n},\varrho)\|&=\|\mathcal{R}f,B^{s,\varphi}_{p}(\mathbb{R}^{n})\|+\varrho^{s}\varphi(\varrho)\|\mathcal{R}f,B^{0}_{p}(\mathbb{R}^{n})\|\\ &\leq c_{1}(\|f,L_{p}(\mathbb{R}^{n},l^{s,\varphi}_{p})\|+\varrho^{s}\varphi(\varrho)\|f,L_{p}(\mathbb{R}^{n},l_{p})\|),\\ \end{split}$ (4.9)
@@
$\begin{split}&\|f,L_{p}(\mathbb{R}^{n},l^{s,\varphi}_{p})\|^{p}+(\varrho^{s}\varphi(\varrho)\|f,L_{p}(\mathbb{R}^{n},l_{p}\|)^{p}\\ =&\int\limits_{\mathbb{R}^{n}}\|f(x),l^{s,\varphi}_{p}\|^{p}dx+(\varrho^{s}\varphi(\varrho))^{p}\int\limits_{\mathbb{R}^{n}}\|f(x),l_{p}\|^{p}dx\\ =&\int\limits_{\mathbb{R}^{n}}\sum^{\infty}_{k=0}(2^{sk}\varphi(2^{k})|f_{k}(x)|)^{p}dx+(\rho^{s}\varphi(\varrho))^{p}\int\limits_{\mathbb{R}^{n}}\sum^{\infty}_{k=0}|f_{k}(x)|^{p}dx\\ =&\int\limits_{\mathbb{R}^{n}}\sum^{\infty}_{k=0}\varkappa^{p}_{k}(s,\varphi,p,\varrho)|f_{k}(x)|^{p}dx=\|f,L_{p}(\mathbb{R}^{n},l^{s,\varphi}_{p}(\varrho))\|^{p},\\ \end{split}$ (4.10)
@@
because the functor $[\cdot,\cdot]_{\theta}$ is exact of type $\theta$ (i.e. it satisfies (A.2); see Appendix A).
@@
$[l^{s_{0},\varphi_{0}}_{p_{0}}(\varrho),l^{s_{1},\varphi_{1}}_{p_{1}}(\varrho)]_{\theta}=l^{s,\varphi}_{p}(\varrho),$ (4.22)
@@
The other cases for $\Omega$ and $G$ are deduced from the case just considered by the same reasoning as that used in [44, Proof of Lemma 4.1] for quadratic interpolation between parameter-dependent Sobolev spaces with the integral-exponent $p=2$. ∎
@@
where $s\in\mathbb{R}$, $p\in(1,\infty)$, and $r:=\max\{2q,m+1\}$. Here, $\widetilde{H}^{\sigma,p,(\mu)}(\Omega,\varrho)$ stands for the Roitberg–Sobolev space with the smoothness index $\sigma\in\mathbb{R}$, integral-exponent $p\in(1,\infty)$, specific integer-valued index $\mu\geq 0$, the norm in this space depending on the parameter $\varrho>0$. If $\sigma\geq\max\{\mu-1+1/p,0\}$, then $\widetilde{H}^{\sigma,p,(\mu)}(\Omega,\varrho)=H^{\sigma}_{p}(\Omega,\varrho)$ with the equivalence of norms in which constants do not depend on $\varrho\geq 1$ [55, Section 9.1.2, p. 298]. Roitberg [55, Section 1.13] defines parameter-dependent norms in $H^{\sigma}_{p}(\Omega)$ and $B^{\sigma}_{p}(\Gamma)$ in a different way then that we use; however, these norms are equivalent to ours with constants not depending on the parameter provided that $\sigma>0$ (see Remark 4.4 given below). Thus, the above-mentioned result [55, Theorem 9.1.1] contains assertion (ii) in the $\varphi(\cdot)\equiv 1$ case because $s>\max\{2q,m+1/p\}$ by the hypothesis.
@@
If $p\neq 2$, then assertion (ii) for every $\varphi\in\Upsilon$ is deduced on the base of Lemma 4.2. Consider first the case where $p\neq 2$ and $2q>m+1/p$. Let us choose a number $p_{1}$ such that $21<p_{1}<p<$ 2 or $2<p<p_{1}<\infty$ and define parameters $\theta\in(0,1)$ and $\varphi_{0}\in\Upsilon$ according to Theorem 2.3. If $21<p_{1}<p<$ 2, we choose a number $s_{1}$ such that $2q<s_{1}<s$ and then define a number $s_{0}>s$ by the formula $s=(1-\theta)s_{0}+\theta s_{1}$. If $2<p<p_{1}$, we choose a number $s_{0}$ such that $2q<s_{0}<s$ and then define a number $s_{1}>s$ by the same formula. By assertion (i), we have the isomorphisms
@@
for every $\lambda\in K$ with $|\lambda|\geq\max\{\lambda_{0},1\}$. Since the interpolation functor $[\cdot,\cdot]_{\theta}$ is exact of type $\theta$, the norms of operator (4.27) and its inverse are bounded from above by the same number $c$. It follows from this by Lemma 4.2 that the two-sided estimate (4.4) holds true with constants that do not depend on $\lambda$.
@@
We recall the definition and some properties of the quadratic interpolation with function parameter between Hilbert spaces. This interpolation is systematically used in the paper together with the well-known complex interpolation (with number parameter) between Banach spaces. Such a quadratic interpolation was introduced by Foiaş and Lions [15, p. 278]. Expounding it, we mainly follow monograph [44, Sect. 1.1]. It is sufficient for our purposes to restrict ourselves to separable Hilbert spaces.
Let $E_{0}$ and $E_{1}$ be separable complex Hilbert spaces such that $E_{1}$ is a dense manifold in $E_{0}$ and that $\|w,E_{0}\|\leq c\,\|w,E_{1}\|$ for every vector $w\in E_{1}$, with the number $c>0$ not depending on $w$ (in other words, the continuous dense embedding $E_{1}\hookrightarrow E_{0}$ holds true). The ordered pair $(E_{0},E_{1})$ of such spaces is called regular. For it, there exists a unique positive-definite self-adjoint operator $J$ given in the Hilbert space $E_{0}$ and such that $E_{1}$ is the domain of $J$ and that $\|Jw,E_{0}\|=\|w,E_{1}\|$ for every $w\in E_{1}$.
Let $\mathcal{B}$ denote the set of all Borel measurable functions $\psi:(0,\infty)\rightarrow(0,\infty)$ such that $\psi$ is bounded on each compact interval $[a,b]$, with $0<a<b<\infty$, and that $1/\psi$ is bounded on every set $[r,\infty)$, with $r>0$. Given $\psi\in\mathcal{B}$ and applying Spectral Theorem to the self-adjoint operator $J$, we obtain the (generally, unbounded) operator $\psi(J)$ in $E_{0}$. Let $(E_{0},E_{1})_{\psi}$ denote the domain of $\psi(J)$ endowed with the norm $\|u,(E_{0},E_{1})_{\psi}\|:=\|\psi(J)u,E_{0}\|$. The space $(E_{0},E_{1})_{\psi}$ is Hilbert and separable with respect to this norm and is continuously embedded in $E_{0}$.
@@
the restriction of T to Ej is a bounded operator T:EjGj for each $\mbox{the restriction of $T$ to $E_{j}$ is a bounded operator $T:E_{j}\to G_{j}$ for each $j\in\{0,1\}$},$ (A.1)
then the restriction of $T$ to $(E_{0},E_{1})_{\psi}$ is also a bounded operator from $(E_{0},E_{1})_{\psi}$ to $(G_{0},G_{1})_{\psi}$. We say in this case that $(E_{0},E_{1})_{\psi}$ is obtained by the quadratic interpolation with the function parameter $\psi$ between the spaces $E_{0}$ and $E_{1}$ and that $T:(E_{0},E_{1})_{\psi}\to(G_{0},G_{1})_{\psi}$ is the result of this interpolation applied to the operators $T:E_{j}\to G_{j}$ with $j\in\{0,1\}$.
A function $\psi\in\mathcal{B}$ is an interpolation parameter if and only if $\psi$ is pseudoconcave on the set $(r,\infty)$ for certain $r>0$, i.e. there exists a positive concave function $\psi_{1}(t)$ of $t\gg 1$ such that the functions $\psi/\psi_{1}$ and $\psi_{1}/\psi$ are bounded on this set. This fundamental fact follows from Peetre’s [54] description of all interpolation functions of positive order [44, Theorem 1.9]. Specifically, the function $\psi(t):=t^{\sigma}\psi_{0}(t)$ of $t\geq 1$ is an interpolation parameter provided that the number $\sigma$ satisfies $10<\sigma<$ 1 and that the function $\psi_{0}:[1,\infty)\to(0,\infty)$ is Borel measurable and varies slowly at infinity (in the sense of Karamata) [44, Theorem 1.11].
@@
Here, $(E_{0},E_{1})$ and $(G_{0},G_{1})$ are arbitrary interpolation pairs of Banach spaces, whereas $T$ is an arbitrary linear mapping acting from whole $E_{0}+E_{1}$ to $G_{0}+G_{1}$ and satisfying (A.1). This estimate means that the interpolation functor $[\cdot,\cdot]_{\theta}$ is exact of type $\theta$ (see, e.g., [62, Theorem 1.9.3(a) and Definition 1.2.2/2]). This property is also used in the proof of Lemma 4.2.

### v2

@@ 3. Applications to elliptic problems
Assertion (i) in the $\ell=0$ case follows from [31, Theorem 5.3] with regard to [9, Theorem 3.4] (see also [29, Theorem 1], where the proof is given). The $\ell\geq 1$ case is reduced to the previous one because the differential operator $\partial/\partial x_{j}$ acts continuously from $H^{s,\varphi}_{p}(\mathbb{R}^{n})$ to $H^{s-1,\varphi}_{p}(\mathbb{R}^{n})$ for every $s\in\mathbb{R}$. Assertion (ii) is proved by a modification of the reasoning given in [64, Proof of Theorem 13.2] in the $\ell=0$ case. Namely, we may assume that $0\in V$; then (3.22) implies the continuous embedding $H^{\ell+n/p,\varphi}_{p}(V_{0})\hookrightarrow C^{\ell}(\overline{V_{0}})$ for a certain ball $V_{0}$ centered at 0, which gives the estimate
@@
for every $w\in\mathcal{S}(\mathbb{R}^{n})$ and each partial derivative $\partial^{\alpha}$ of the order $|\alpha|\leq\ell$. Hence, the delta-function $\delta$ and all its partial derivatives $\partial^{\alpha}\delta$ with $|\alpha|\leq\ell$ belong to the dual of $H^{\ell+n/p,\varphi}_{p}(\mathbb{R}^{n})$. Since this dual equals $H^{-\ell-n/p,1/\varphi}_{p^{\prime}}(\mathbb{R}^{n})$ [64, Section 13, Subsection 3], we have the inclusion $\partial^{\alpha}\Phi\delta=\Phi(\partial^{\alpha}\delta)\in H^{-\ell-n/p}_{p^{\prime}}(\mathbb{R}^{n})$ whatever $0\leq|\alpha|\leq\ell$, with $\Phi w:=\mathcal{F}^{-1}[\varphi^{-1}(\langle\xi\rangle)\mathcal{F}w(\xi)]$ whenever $w\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$. Hence, $\Phi\delta\in H^{-n/p}_{p^{\prime}}(\mathbb{R}^{n})$ (see, e.g., [61, Theorem 2.3.8(ii)]), which gives $\delta\in H^{-n/p,1/\varphi}_{p^{\prime}}(\mathbb{R}^{n})$. The last inclusion means that the function $\mu(\xi):=\langle\xi\rangle^{n/p}\varphi(\langle\xi\rangle)$ of $\xi\in\mathbb{R}^{n}$ satisfies $\mathcal{F}^{-1}\mu^{-1}(\xi)\in L_{p^{\prime}}(\mathbb{R}^{n})$. This inclusion is equivalent to (2.32) due to [64, Theorem 13.2] and [31, Theorem 5.3]. Thus, (3.22) implies (2.32).
@@ 4. Applications to parameter-elliptic boundary-value problems
    For arbitrarily fixed point $x\in\Gamma$, vector $\xi\in\penalty 10000\mathbb{R}^{n}$, tangent to the boundary $\Gamma$ at the point $x$, and the parameter $\lambda\in K$ such that $|\xi|+|\lambda|\neq 0$, the polynomials $B^{\circ}_{j}(x;\xi+\tau\nu(x),\lambda)$, $j=1,\ldots,q$, in $\tau\in\mathbb{C}$ are linearly independent modulo the polynomial $\prod_{j=1}^{q}(\tau-\tau^{+}_{j}(x;\xi,\lambda))$. Here, $\tau^{+}_{1}(x;\xi,\lambda),\ldots,\tau^{+}_{q}(x;\xi,\lambda)$ are all $\tau$-roots of the polynomial $A^{\circ}(x;\xi+\tau\nu(x),\lambda)$ that have the positive imaginary part and are written with regard for their multiplicity, whereas $\nu(x)$ stands for the unit vector of the inner normal to $\Gamma$ at $x$.
We have used the following notations:
@@
    There exists a number $\lambda_{0}>0$ such that, for each $\lambda\in K$ with $|\lambda|\geq\penalty 10000\lambda_{0}$, operator (4.3) is an isomorphism of $H^{s,\varphi}_{p}(\Omega)$ onto $\Xi^{s,\varphi}_{p}(\Omega,\Gamma)$ whatever $p\in(1,\infty)$, $s>m+1/p$, and $\varphi\in\Upsilon$.
@@
    $\quad+\sum_{j=1}^{q}\,\bigl(\|B_{j}(\lambda)u,B^{s-m_{j}-1/p,\varphi}_{p}(\Gamma)\|+|\lambda|^{s-m_{j}-1/p}\varphi(|\lambda|)\,\|B_{j}(\lambda)u,B^{0}_{p}(\Gamma)\|\bigr)$
@@
Note that $L_{p}(\Omega)=H^{0}_{p}(\Omega)$ with equality of norms. Since the ellipticity of the parameter-dependent problem (4.1), (4.2) in $K$ implies its ellipticity in the angle $K:=\{-\lambda:\lambda\in K\}$, the conclusions of this theorem is also valid for $\lambda\in-K$ with $|\lambda|\gg 1$.
@@
[…]
@@
[…]
@@
because the functor $\cdot,\cdot]_{\theta}$ is exact of type $\theta$ (i.e. it satisfies (A.2); see Appendix A).
@@
$l^{s_{0},\varphi_{0}}_{p_{0}}(\varrho),l^{s_{1},\varphi_{1}}_{p_{1}}(\varrho)]_{\theta}=l^{s,\varphi}_{p}(\varrho),$ (4.22)
@@
The other cases for $\Omega$ and $G$ are deduced from the case just considered by the same reasoning as that used in [44, Proof of Lemma 4.1] for quadratic interpolation between parameter-dependent Sobolev spaces with the integrability exponent $p=2$. ∎
@@
where $s\in\mathbb{R}$, $p\in(1,\infty)$, and $r:=\max\{2q,m+1\}$. Here, $\widetilde{H}^{\sigma,p,(\mu)}(\Omega,\varrho)$ stands for the Roitberg–Sobolev space with the smoothness index $\sigma\in\mathbb{R}$, integrability exponent $p\in(1,\infty)$, specific integer-valued index $\mu\geq 0$, the norm in this space depending on the parameter $\varrho>0$. If $\sigma\geq\max\{\mu-1+1/p,0\}$, then $\widetilde{H}^{\sigma,p,(\mu)}(\Omega,\varrho)=H^{\sigma}_{p}(\Omega,\varrho)$ with the equivalence of norms in which constants do not depend on $\varrho\geq 1$ [55, Section 9.1.2, p. 298]. Roitberg [55, Section 1.13] defines parameter-dependent norms in $H^{\sigma}_{p}(\Omega)$ and $B^{\sigma}_{p}(\Gamma)$ in a different way then that we use; however, these norms are equivalent to ours with constants not depending on the parameter provided that $\sigma>0$ (see Remark 4.4 given below). Thus, the above-mentioned result [55, Theorem 9.1.1] contains assertion (ii) in the $\varphi(\cdot)\equiv 1$ case because $s>\max\{2q,m+1/p\}$ by the hypothesis.
@@
If $p\neq 2$, then assertion (ii) for every $\varphi\in\Upsilon$ is deduced on the base of Lemma 4.2. Consider first the case where $p\neq 2$ and $2q>m+1/p$. Let us choose a number $p_{1}$ such that $1<p_{1}<p<2$ or $2<p<p_{1}<\infty$ and define parameters $\theta\in(0,1)$ and $\varphi_{0}\in\Upsilon$ according to Theorem 2.3. If $1<p_{1}<p<2$, we choose a number $s_{1}$ such that $2q<s_{1}<s$ and then define a number $s_{0}>s$ by the formula $s=(1-\theta)s_{0}+\theta s_{1}$. If $2<p<p_{1}$, we choose a number $s_{0}$ such that $2q<s_{0}<s$ and then define a number $s_{1}>s$ by the same formula. By assertion (i), we have the isomorphisms
@@
for every $\lambda\in K$ with $|\lambda|\geq\max\{\lambda_{0},1\}$. Since the interpolation functor $\cdot,\cdot]_{\theta}$ is exact of type $\theta$, the norms of operator (4.27) and its inverse are bounded from above by the same number $c$. It follows from this by Lemma 4.2 that the two-sided estimate (4.4) holds true with constants that do not depend on $\lambda$.
@@
We recall the definition and some properties of the quadratic interpolation with function parameter between Hilbert spaces. This interpolation is systematically used in the paper together with the well-known complex interpolation (with number parameter) between Banach spaces. Such a quadratic interpolation was introduced by Foiaş and Lions [15, p. 278]. Expounding it, we mainly follow the monograph [44, Section 1.1]. It is sufficient for our purposes to restrict ourselves to separable Hilbert spaces.
Let $E_{0}$ and $E_{1}$ be separable complex Hilbert spaces such that $E_{1}$ is a dense linear manifold in $E_{0}$ and that $\|w,E_{0}\|\leq c\,\|w,E_{1}\|$ for every vector $w\in E_{1}$, with the number $c>0$ not depending on $w$ (in other words, the continuous dense embedding $E_{1}\hookrightarrow E_{0}$ holds true). The ordered pair $(E_{0},E_{1})$ of such spaces is called regular. For it, there exists a unique positive-definite self-adjoint operator $J$ given in the Hilbert space $E_{0}$ and such that $E_{1}$ is the domain of $J$ and that $\|Jw,E_{0}\|=\|w,E_{1}\|$ for every $w\in E_{1}$.
Let $\mathcal{B}$ denote the set of all Borel measurable functions $\psi:(0,\infty)\rightarrow(0,\infty)$ such that $\psi$ is bounded on each compact interval $[a,b]$, with $0<a<b<\infty$, and that $1/\psi$ is bounded on every set $r,\infty)$, with $r>0$. Given $\psi\in\mathcal{B}$ and applying the spectral theorem to the self-adjoint operator $J$, we obtain the (generally, unbounded) operator $\psi(J)$ in $E_{0}$. Let $(E_{0},E_{1})_{\psi}$ denote the domain of $\psi(J)$ endowed with the norm $\|u,(E_{0},E_{1})_{\psi}\|:=\|\psi(J)u,E_{0}\|$. The space $(E_{0},E_{1})_{\psi}$ is Hilbert and separable with respect to this norm and is continuously embedded in $E_{0}$.
@@
[…]
then the restriction of $T$ to $(E_{0},E_{1})_{\psi}$ is a bounded operator from $(E_{0},E_{1})_{\psi}$ to $(G_{0},G_{1})_{\psi}$. We say in this case that $(E_{0},E_{1})_{\psi}$ is obtained by the quadratic interpolation with the function parameter $\psi$ between the spaces $E_{0}$ and $E_{1}$ and that $T:(E_{0},E_{1})_{\psi}\to(G_{0},G_{1})_{\psi}$ is the result of this interpolation applied to the operators $T:E_{j}\to G_{j}$ with $j\in\{0,1\}$.
A function $\psi\in\mathcal{B}$ is an interpolation parameter if and only if $\psi$ is pseudoconcave on the set $(r,\infty)$ for certain $r>0$, i.e. there exists a positive concave function $\psi_{1}(t)$ of $t\gg 1$ such that the functions $\psi/\psi_{1}$ and $\psi_{1}/\psi$ are bounded on this set. This fundamental fact follows from Peetre’s [54] description of all interpolation functions of positive order [44, Theorem 1.9]. Specifically, the function $\psi(t):=t^{\sigma}\psi_{0}(t)$ of $t\geq 1$ is an interpolation parameter provided that the number $\sigma$ satisfies $0<\sigma<1$ and that the function $\psi_{0}:[1,\infty)\to(0,\infty)$ is Borel measurable and varies slowly at infinity (in the sense of Karamata) [44, Theorem 1.11].
@@
Here, $(E_{0},E_{1})$ and $(G_{0},G_{1})$ are arbitrary interpolation pairs of Banach spaces, whereas $T$ is an arbitrary linear mapping acting from whole $E_{0}+E_{1}$ to $G_{0}+G_{1}$ and satisfying (A.1). This estimate means that the interpolation functor $\cdot,\cdot]_{\theta}$ is exact of type $\theta$ (see, e.g., [62, Theorem 1.9.3(a) and Definition 1.2.2/2]). This property is also used in the proof of Lemma 4.2.